OpenStudy (anonymous):

(PLEASE HELP) At a flower shop, Betsy paid $33 for 3 roses and 9 tulips. Bert paid $48 for 12 roses and 8 tulips. Find the cost of one rose and one tulip without using systems of equations

OpenStudy (e.cociuba):

For equation r=Rose t=Tulip. Do 33=3r+9t and 48=12r+8t and you do 4(33r+9t) and that is 123=12r+36t, than you do 123=12r+36t-48=12r+8t and that is 84=28t and that is t=3. Than I plugged 3 into 33=3r+9(3) and 33=3r+27, 3r=6 r=2. For non equation I plugged in random prices to 3 roses and 9 tulips to get to 33. 3X2=6 and 9X2=18 and that is 24. and 3X3=9 and 9X3-27 and that equals 36, that too much so I did 3X2=6 and 9X3-27 and that equals 33.

OpenStudy (anonymous):

Thank you

OpenStudy (e.cociuba):

Np:)